# Kernel Mean Embedding Topology: Weak and Strong Forms for Stochastic Kernels and Implications for Model Learning

### v1

@@ 3.3 Comparison of Weak Topologies
$I^{-1}\circ\gamma_{\theta}(y)$ converges vaguely to $I^{-1}\circ\gamma(y)$. Note that $\{I^{-1}\circ\gamma_{\theta}(y)\}\subset{\mathcal{G}}$ and ${\mathcal{G}}$ is tight, and so, $\{I^{-1}\circ\gamma_{\theta}(y)\}$ has a subsequence which converges to some probability measure in weak convergence topology (and also in vague topology) by Prohorov’s theorem [30, Theorem 1.4.12]. This implies that $I^{-1}\circ\gamma(y)$ is a probability measure. Hence, $\gamma\in L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})$. Therefore, the set $\mathop{\rm conv}L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ has the closure with respect to the norm-topology (or equivalently, with respect to the weak kernel mean embedding topology) in $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})\equiv\Gamma$. This implies that $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ has the closure with respect to the weak kernel mean embedding topology in $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})\equiv\Gamma$. Conversely, suppose that the set $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ has the closure with respect to the weak kernel mean embedding topology in $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})\equiv\Gamma$ and ${\mathsf{U}}$ is complete. Let $\{\nu_{n}\}$ be a sequence in ${\mathcal{G}}$ and for each $n$, define the constant function $\gamma_{n}(y)=I\circ\nu_{n}$ for all $y\in{\mathsf{Y}}$. Since $\mathcal{L}_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ is relatively (sequentially) compact with respect to the weak kernel mean embedding topology, there exists a subsequence $\{\gamma_{n_{k}}\}$ of $\{\gamma_{n}\}$ that converges with respect to the weak kernel mean embedding topology. This implies that subsequence $\{\nu_{n_{k}}\}$ of $\{\nu_{n}\}$ also converges with respect to the weak MMD topology. According to [55, Lemma 5], we also have convergence in the weak convergence topology, as the limit measure of ${\nu_{n_{k}}}$ is a probability measure, given that the closure of $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ lies within $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})$. Hence, ${\mathcal{G}}$ is relatively (sequentially) compact with respect to the weak convergence topology. By Prohorov’s theorem [30, Theorem 1.4.12], ${\mathcal{G}}$ is tight. This completes the proof of item (ii). ∎
@@ 4 Strong Kernel Mean Embedding Topology for Stochastic Kernels
Weak topologies are primarily introduced to prove the existence of optimal policies in decision-making problems under the most general and mild assumptions. Their mathematical structure is well-suited for demonstrating such existence results, as weak topologies simplify the handling of compactness and continuity properties in infinite-dimensional spaces. Consequently, it is both natural and effective to define weak topologies on the set of policies. For instance, the Young narrow topology was originally defined to establish the existence of optimal relaxed policies in continuous-time deterministic optimal control problems [64]. Similarly, the $w^{*}$-topology on the set of stochastic kernels was developed in [11, Section 2.4] to analyze randomized Markov policies and prove the existence of optimal policies for continuous-time stochastic control problems under the average cost optimality criterion. More recently, in [51], the $w^{*}$-topology was employed to establish the existence of optimal policies in team decision problems (see also [68, 53]). We note that the weak kernel mean embedding topology is equivalent to both the $w^{*}$-topology and the Young narrow topology for proving the existence of optimal solutions, as established in Theorem 3.10 and Theorem 3.11. Consequently, all results demonstrated in the aforementioned papers also hold under the weak kernel mean embedding topology.
@@
The weak convergence topology on the set of probability measures can be metrized using various metrics, one of which is the bounded-Lipschitz metric, denoted by $\rho_{BL}$ [20, p. 394]. Using this metric, we can define the following strong topology, similar to the strong kernel mean embedding topology, on the set of stochastic kernels $\Gamma$: $\gamma_{\lambda}\rightarrow\gamma$ in $\Gamma$ if
@@
Then, $\gamma_{\lambda}\rightarrow\gamma$ in strong kernel mean embedding topology.
@@
Suppose that the result is not true. Then, there exists a subsequence $\{\gamma_{\tau}\}$ of $\{\gamma_{\lambda}\}$ such that the kernel mean embedding strong norm of the difference $\gamma_{\tau}-\gamma$ is strictly greater than 00 for all $\tau$. Under the stated assumptions and by Arzela-Ascoli theorem [50, Theorem 4.44], there exists a subsequence $\{f_{\theta}\}$ of $\{f_{\tau}\}$ such that $f_{\theta}\rightarrow\tilde{f}$ uniformly over compact subsets, for some continuous and bounded density function $\tilde{f}$. It is then very simple to establish that $\gamma_{\theta}\rightarrow\tilde{\gamma}$ with respect to MMD topology uniformly over compact subsets, where $\tilde{\gamma}(y)(du)\triangleq\tilde{f}(u,y)\,\xi_{{\mathsf{abs}}}(du)$. Since $\gamma_{n}\rightharpoonup^{*}\gamma$, we can also conclude that $\gamma=\tilde{\gamma}$ $\mu$-almost-surely. As uniform convergence over compact sets is stronger than pointwise convergence, we have
[…]
This implies the convergence in strong kernel mean embedding topology of $\gamma_{\theta}$ to $\gamma$. This contradicts with the initial assumption, which completes the proof. ∎
@@
Although condition (b) in the theorem is not overly restrictive, as explained in Remark 9, it can be slightly relaxed using the Arzelà-Ascoli theorem for vector-valued functions [46, Theorem 1.6.14]. Specifically, we aim to show that the mappings $\gamma_{\lambda}:{\mathsf{Y}}\rightarrow{\mathcal{H}}_{{\mathcal{P}}}$ are relatively sequentially compact under the topology of uniform convergence on compacta. To achieve this, two requirements must be met:
@@
Since the MMD topology is equivalent to the weak convergence topology on ${\mathcal{H}}_{{\mathcal{P}}}$ [55, Lemma 5 and Remark 6], the latter condition is equivalent to the tightness of $\{\gamma_{\lambda}(y)(\cdot)\}$ by Prokhorov’s theorem [30, Theorem 1.4.12]. Consequently, condition (b) in the theorem can be replaced with the following slightly relaxed version:
@@ 5 Applications
5.1 Robustness of Markov Decision Processes
@@
    ${\mathsf{X}}$ is $\sigma$-compact locally compact metric space and ${\mathsf{A}}$ is finite.
Note that if ${\mathsf{X}}=\mathbb{R}^{n}$ for some $n\geq 1$, then ${\mathsf{X}}$ is $\sigma$-compact locally compact metric space under Euclidean or one of the equivalent metrics. Moreover, by [54, Theorem 3.2], under the weak continuity of the transition kernel $p$ (i.e., $p(\cdot|x_{n},a_{n})\rightarrow p(\cdot|x,a)$ in weak convergence topology for any $(x_{n},a_{n})\rightarrow(x,a)$ in ${\mathsf{X}}\times{\mathsf{A}}$), any MDP with compact action space can be approximated by MDPs with finite action spaces. Therefore, finite action space assumption is not restrictive in terms of optimality.
@@ 5.1.1 Robustness Under Strong Kernel Mean Embedding Topology
that is, $p_{n}$ converges to $p$ uniformly with respect to MMD topology. Since MMD topology is equivalent to weak convergence topology under Assumption 1 [55, Lemma 5 and Remark 6], we have uniform convergence with respect to the weak convergence topology. But this uniform convergence implies the following:
@@
[32, Theorem 4.2, Theorem 4.4] Suppose that $\{p_{n}\}$ and $p$ are weakly continuous and $p_{n}(\cdot|x_{n},a_{n})\rightarrow p(\cdot|x,a)\,\,\text{weakly whenever}\,\,(x_{n},a_{n})\rightarrow(x,a)$. Then, we have
@@ 5.1.2 Robustness Under Weak Kernel Mean Embedding Topology
Under Assumption 4 and by Arzela-Ascoli theorem [50, Theorem 4.44], there exists a subsequence $\{f_{\theta}\}$ of $\{f_{\lambda}\}$ such that $f_{\theta}\rightarrow\tilde{f}$ uniformly over compact subsets, for some continuous and bounded density function $\tilde{f}$. It is then very simple to establish that $p_{\theta}\rightarrow\tilde{p}$ with respect to MMD topology uniformly over compact subsets, where $\tilde{p}(dy|x,a)\triangleq\tilde{f}(y|x,a)\,\kappa_{{\mathsf{abs}}}(dy)$. Since $p_{n}\rightharpoonup^{*}p$, we can also conclude that $p=\tilde{p}$ $\mu_{{\mathsf{ref}}}$-almost-surely. As uniform convergence over compact sets is stronger than pointwise convergence, we have
[…]
@@
By Assumption 4 and by Arzela-Ascoli theorem [50, Theorem 4.44], there exists a subsequence $\{f_{\alpha}(y|\cdot,\cdot)\}$ of $\{f_{\xi}(y|\cdot,\cdot)\}$ such that $f_{\alpha}(y|\cdot,\cdot)\rightarrow\tilde{f}(y|\cdot,\cdot)$ uniformly over compact subsets, for some continuous and bounded density function $\tilde{f}(y|\cdot,\cdot)$. Define the compact set $K\triangleq\{(x_{\alpha},a_{\alpha})\}\bigcup\{(x,a)\}$. Then we have
@@
This is contradicting with
@@ 5.1.3 Implications for Asymptotic Optimality under Empirical Model Learning
The implications of the results given above is that one can learn models via empirical data. In particular, there exist several studies which ensures that kernels can be learned from data, see e.g. [28, Theorem 4] for one of the earlier results under a stronger convergence criterion in the presence of empirical data; and [62] and the references therein for an empirical convergence analysis under the strong kernel mean embedding topology.
For an MDP with Borel state and action spaces, under weak continuity conditions, it is known that the action space can be approximated by a finite action set with arbitrarily small change in performance; see [54] [52, Theorem 3.16] for the discounted cost and [54],[52, Theorem 3.22] for average cost criterion. Thus, one needs to learn only finitely many kernels from data. Given the analysis presented in the previous section, these then lead to robustness to empirical learning in MDPs; see [44] [32, 31] for an application.

### v2

@@ 3.3 Comparison of Weak Topologies
$I^{-1}\circ\gamma_{\theta}(y)$ converges vaguely to $I^{-1}\circ\gamma(y)$. Note that $\{I^{-1}\circ\gamma_{\theta}(y)\}\subset{\mathcal{G}}$ and $\mathcal{G}$} is tight, and so, $\{I^{-1}\circ\gamma_{\theta}(y)\}$ has a subsequence which converges to some probability measure in weak convergence topology (and also in vague topology) by Prohorov’s theorem [39, Theorem 1.4.12]. This implies that $I^{-1}\circ\gamma(y)$ is a probability measure. Hence, $\gamma\in L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})$. Therefore, the set $\mathop{\rm conv}L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ has the closure with respect to the norm-topology (or equivalently, with respect to the weak kernel mean embedding topology) in $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})\equiv\Gamma$. This implies that $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ has the closure with respect to the weak kernel mean embedding topology in $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})\equiv\Gamma$. Conversely, suppose that the set $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ has the closure with respect to the weak kernel mean embedding topology in $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})\equiv\Gamma$ and $\mathsf{U}$} is complete. Let $\{\nu_{n}\}$ be a sequence in $\mathcal{G}$} and for each $n$, define the constant function $\gamma_{n}(y)=I\circ\nu_{n}$ for all $y\in{\mathsf{Y}}$. Since $\mathcal{L}_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ is relatively (sequentially) compact with respect to the weak kernel mean embedding topology, there exists a subsequence $\{\gamma_{n_{k}}\}$ of $\{\gamma_{n}\}$ that converges with respect to the weak kernel mean embedding topology. This implies that subsequence $\{\nu_{n_{k}}\}$ of $\{\nu_{n}\}$ also converges with respect to the weak MMD topology. According to [67, Lemma 5], we also have convergence in the weak convergence topology, as the limit measure of $\nu_{n_{k}}$} is a probability measure, given that the closure of $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ lies within $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})$. Hence, $\mathcal{G}$} is relatively (sequentially) compact with respect to the weak convergence topology. By Prohorov’s theorem [39, Theorem 1.4.12], $\mathcal{G}$} is tight. This completes the proof of item (ii). ∎
@@ 4 Strong Kernel Mean Embedding Topology for Stochastic Kernels
Weak topologies are primarily introduced to prove the existence of optimal policies in decision-making problems under the most general and mild assumptions. Their mathematical structure is well-suited for demonstrating such existence results, as weak topologies simplify the handling of compactness and continuity properties in infinite-dimensional spaces. Consequently, it is both natural and effective to define weak topologies on the set of policies. For instance, the Young narrow topology was originally defined to establish the existence of optimal relaxed policies in continuous-time deterministic optimal control problems [75]. Similarly, the $w^{*}$-topology on the set of stochastic kernels was developed in [19, Section 2.4] to analyze randomized Markov policies and prove the existence of optimal policies for continuous-time stochastic control problems under the average cost optimality criterion. More recently, in [63], the $w^{*}$-topology was employed to establish the existence of optimal policies in team decision problems (see also [79, 65]). We note that the weak kernel mean embedding topology is equivalent to both the $w^{*}$-topology and the Young narrow topology for proving the existence of optimal solutions, as established in Theorem 3.10 and Theorem 3.11. Consequently, all results demonstrated in the aforementioned papers also hold under the weak kernel mean embedding topology.
@@
The weak convergence topology on the set of probability measures can be metrized using various metrics, one of which is the bounded-Lipschitz metric, denoted by $\rho_{BL}$ [28, p. 394]. Using this metric, we can define the following strong topology, similar to the strong kernel mean embedding topology, on the set of stochastic kernels $\Gamma$: $\gamma_{\lambda}\rightarrow\gamma$ in $\Gamma$ if
@@
Then, $\gamma_{\lambda}\rightarrow\gamma$ in strong kernel mean embedding topology for $q<\infty$.
@@
Suppose that the result is not true. Then, there exists a subsequence $\{\gamma_{\tau}\}$ of $\{\gamma_{\lambda}\}$ such that the kernel mean embedding strong norm of the difference $\gamma_{\tau}-\gamma$ is strictly greater than some $\epsilon>0$ for all $\tau$. Under the stated assumptions and by Arzela-Ascoli theorem [62, Theorem 4.44], there exists a subsequence $\{f_{\theta}\}$ of $\{f_{\tau}\}$ such that $f_{\theta}\rightarrow\tilde{f}$ uniformly over compact subsets, for some continuous and bounded density function $\tilde{f}$. It then follows that $\gamma_{\theta}\rightarrow\tilde{\gamma}$ with respect to MMD topology uniformly over compact subsets, where $\tilde{\gamma}(y)(du)\triangleq\tilde{f}(u,y)\,\xi_{{\mathsf{abs}}}(du)$. Since $\gamma_{n}\rightharpoonup^{*}\gamma$, we can also conclude that $\gamma=\tilde{\gamma}$ $\mu$-almost-surely. As uniform convergence over compact sets is stronger than pointwise convergence, we have
$\|\gamma_{\theta}(y)(\cdot)-\gamma(y)(\cdot)\|_{{\mathcal{M}}}\rightarrow 0\,\,\text{$\mu$-almost-surely}.$
This implies the convergence in strong kernel mean embedding topology for $q<\infty$ of $\gamma_{\theta}$ to $\gamma$ as the MMD norm of the difference of two probability measures is bounded via Assumption 1-(a). This contradicts with the initial assumption, which completes the proof. ∎
@@
Although condition (b) in the theorem is not overly restrictive, as explained in Remark 10, it can be slightly relaxed using the Arzelà-Ascoli theorem for vector-valued functions [56, Theorem 1.6.14]. Specifically, we aim to show that the mappings $\gamma_{\lambda}:{\mathsf{Y}}\rightarrow{\mathcal{H}}_{{\mathcal{P}}}$ are relatively sequentially compact under the topology of uniform convergence on compacta. To achieve this, two requirements must be met:
@@
Since the MMD topology is equivalent to the weak convergence topology on ${\mathcal{H}}_{{\mathcal{P}}}$ [67, Lemma 5 and Remark 6], the latter condition is equivalent to the tightness of $\{\gamma_{\lambda}(y)(\cdot)\}$ by Prokhorov’s theorem [39, Theorem 1.4.12]. Consequently, condition (b) in the theorem can be replaced with the following slightly relaxed version:
@@
Remark 9.
The result in Theorem 4.2 can alternatively be established under a more direct assumption:
$f_{\lambda}(\cdot,y)\to f(\cdot,y)\quad\xi_{{\mathsf{abs}}}\text{-a.s.}$
for all $y\in{\mathsf{Y}}$, except on a set of $\mu$-measure zero. In that case, Scheffe’s theorem [17, Theorem 16.12] can be directly applied to obtain the desired conclusion. It is worth emphasizing that, under the assumptions (a) and (b) of Theorem 4.2, the Arzela-Ascoli theorem guarantees uniform convergence of densities over compact subsets along a subsequence. However, such uniform convergence on compact sets does not, in general, imply almost sure pointwise convergence along the original sequence – simple counterexamples can be constructed. In Theorem 4.2, we additionally assume that the stochastic kernels converge in the weak kernel mean embedding topology, which ensures that the corresponding sequence of densities admits a unique cluster point almost surely. Combined with the Arzela-Ascoli theorem, this implies almost sure pointwise convergence, which is precisely the condition required by Scheffe’s theorem. Hence, the approach based directly on Scheffe’s theorem yields a stronger result than ours. Nevertheless, the assumptions in Theorem 4.2 are easier to verify in practice, which motivates stating the weaker, but more accessible, version presented in this paper.
@@ 5 Applications
5.1 Robustness of Markov Decision Processes under Strong Kernel Topologies
@@
    $\mathsf{X}$} is locally compact metric space and $\mathsf{A}$} is finite.
Note that if ${\mathsf{X}}=\mathbb{R}^{n}$ for some $n\geq 1$, then $\mathsf{X}$} is locally compact metric space under Euclidean or one of the equivalent metrics. Moreover, by [66, Theorem 3.2], under the weak continuity of the transition kernel $p$ (i.e., $p(\cdot|x_{n},a_{n})\rightarrow p(\cdot|x,a)$ in weak convergence topology for any $(x_{n},a_{n})\rightarrow(x,a)$ in ${\mathsf{X}}\times{\mathsf{A}}$), any MDP with compact action space can be approximated by MDPs with finite action spaces. Therefore, finite action space assumption is not restrictive in terms of optimality.
@@ 5.1.1 Robustness Under Strong Kernel Mean Embedding Topology
that is, $p_{n}$ converges to $p$ uniformly with respect to MMD topology. Since MMD topology is equivalent to weak convergence topology under Assumption 1 [67, Lemma 5 and Remark 6], we have uniform convergence with respect to the weak convergence topology. But this uniform convergence implies the following:
@@
[41, Theorem 4.2, Theorem 4.4] Suppose that $\{p_{n}\}$ and $p$ are weakly continuous and $p_{n}(\cdot|x_{n},a_{n})\rightarrow p(\cdot|x,a)\,\,\text{weakly whenever}\,\,(x_{n},a_{n})\rightarrow(x,a)$. Then, we have
@@ 5.1.2 Robustness Under Weak Kernel Mean Embedding Topology
Under Assumption 4 and by Arzela-Ascoli theorem [62, Theorem 4.44], there exists a subsequence $\{f_{\theta}\}$ of $\{f_{\lambda}\}$ such that $f_{\theta}\rightarrow\tilde{f}$ uniformly over compact subsets, for some continuous and bounded density function $\tilde{f}$. It is then very simple to establish that $p_{\theta}\rightarrow\tilde{p}$ with respect to MMD topology uniformly over compact subsets, where $\tilde{p}(dy|x,a)\triangleq\tilde{f}(y|x,a)\,\kappa_{{\mathsf{abs}}}(dy)$. Since $p_{n}\rightharpoonup^{*}p$, we can also conclude that $p=\tilde{p}$ $\mu_{{\mathsf{ref}}}$-almost-surely. As uniform convergence over compact sets is stronger than pointwise convergence, we have
$\|p_{\theta}(\cdot|x,a)-p(\cdot|x,a)\|_{{\mathcal{M}}}\rightarrow 0\,\,\text{$\mu_{{\mathsf{ref}}}$-almost-surely}.$
@@
By Assumption 4 and by Arzela-Ascoli theorem [62, Theorem 4.44], there exists a subsequence $\{f_{\alpha}(y|\cdot,\cdot)\}$ of $\{f_{\xi}(y|\cdot,\cdot)\}$ such that $f_{\alpha}(y|\cdot,\cdot)\rightarrow\tilde{f}(y|\cdot,\cdot)$ uniformly over compact subsets, for some continuous and bounded density function $\tilde{f}(y|\cdot,\cdot)$. Define the compact set $K\triangleq\{(x_{\alpha},a_{\alpha})\}\bigcup\{(x,a)\}$. Then we have
@@
This contradicts with
@@
Remark 11.
Theorem 5.2 relies on continuous convergence of densities; that is,
$f_{n}(\cdot|x_{n},a_{n})\rightarrow f(\cdot|x,a),$ (13)
where $(x_{n},a_{n})\rightarrow(x,a)$ (see [53, Proposition 2.8] for a similar result under weak convergence of transition kernels). This is indeed equivalent to the uniform convergence of densities over compact sets when densities are continuous and this is established through uniform boundedness, equicontinuity, and the Arzela-Ascoli theorem along a subsequence in the above proof. However, we note that the same result can alternatively be obtained by directly assuming, instead of Assumption 4, the continuous convergence of densities (13) and applying Scheffe’s theorem [17, Theorem 16.12].
@@ 5.1.3 Implications for Asymptotic Optimality under Empirical Model Learning
The implications of the results given above is that one can learn models via empirical data. In particular, there exist several studies which ensures that kernels can be learned from data, see e.g. [36, Theorem 4] for one of the earlier results under a stronger convergence criterion in the presence of empirical data; and [73] and the references therein for an empirical convergence analysis under the strong kernel mean embedding topology.
For an MDP with Borel state and action spaces, under weak continuity conditions, it is known that the action space can be approximated by a finite action set with arbitrarily small change in performance; see [66] [64, Theorem 3.16] for the discounted cost and [66],[64, Theorem 3.22] for average cost criterion. Thus, one needs to learn only finitely many kernels from data. Given the analysis presented in the previous section, these then lead to robustness to empirical learning in MDPs; see [53, 81] (building on [41, 40]) for an application.
Remark 12 (On Adapted Topologies).
A topology on spaces of probability measures corresponding to laws of stochastic processes, which has been used in a wide variety of contexts in stochastic analysis, is defined by the following. A sequence of stochastic processes is said to converge to another process if their finite-dimensional marginals converge weakly, and their conditional distributions of future variables given the past (viewed as measure-valued stochastic processes) also converge weakly. Aldous has termed this extended weak convergence [1] and Hellwig has named it the information topology [37]; these have recently been shown to be equivalent in discrete-time [5, 55]. The closely related adapted Wasserstein metric [9, 5, 15] has been shown to possess strong robustness properties in a variety of applications [14, 6, 10], not unlike the closely related continuous weak convergence notion we have studied above for MDPs [41] and in controlled diffusion problems [59]. We should note one distinction, however: The convergence of stochastic processes is with regard to the process measure itself and not only on the conditional probability measures at each element in the domain; in particular, a prior measure may entail a more restricted support for future marginal measures and this may affect the convergence properties under the adapted Wasserstein distance whereas the convergence of a stochastic kernel sequence is with regard to the properties of the kernels defined on the entire (domain) space.
5.2 Continuity and Approximations of Control Policies under Weak Kernel Topologies
Under the Young narrow topology or $w^{*}$-topology on control policies, and hence under the weak kernel mean embedding topology (by Theorem 3.10 and Theorem 3.11), one obtains broad continuity and approximation results, as established in several foundational studies. Specifically, [20, 2] showed that, in controlled diffusion processes, the expected cost depends continuously on the control policy, [59, 61, 60] demonstrated that, since piecewise-constant or Lipschitz policies are dense under this topology (see, e.g., [50, 3]), both near-optimality of Lipschitz policies and the validity of discrete-time approximations follow. Related analyses on discrete-time approximations can also be found in [31, 45, 44]. Moreover, invariant measures of diffusions depend continuously on control policies under this topology [2], and an analogous result holds in the discrete-time setting [79]. By the equivalence between the Young topology and the weak kernel mean embedding topology, one can also establish existence and approximation results for systems with decentralized information structures [78, 65, 63]. These findings further extend to mean-field and game-theoretic models, providing desirable compactness, convexity, and continuity properties [49, 7, 13].
In summary, strong kernel mean embedding topology (with the $\infty$-norm) is well-suited for analyzing model convergence and robustness properties, whereas weak kernel mean embedding topology is particularly appropriate for control policy spaces, yielding broad results on continuity, existence, and approximation.